OpenStudy (anonymous):

7.An amusement park charges $8 admission and average of 2000 visitors per day. A survey shows that , for each $1 increase in the admission cost, 100 fewer people would visit the park. b.Find the coordinates of the maximum point of this function

OpenStudy (anonymous):

we need a quadratic function to start

OpenStudy (anonymous):

at present they make \(8\times 2000=\$16000\) what happens if they charge $9 ? they will make $1 more per person, but have 100 fewer people, so they will make \[9\times 1990=\$17910\]

OpenStudy (anonymous):

suppose they charge $10? then they make $2 more per person, but have 200 fewer people, make \[10\times 1980=\$19800\] now lets cut to the chase

OpenStudy (anonymous):

if they charge say \(x\) more dollars they make \(x+8\) per person, but have \(2000-10x\) customers for a total amount of money as \[(x+8)(2000-10x)\] that is your quadratic function maximum is at the vertex

OpenStudy (anonymous):

you good from there?